How do you find f(x+h) for this function: f(x)=x^2-2x+5. I know you pug in (x+h0 for x.. but i don't understand how to solve it. If you could please help me that'd be great. :)
Answer by stanbon(75887)

How do you find f(x+h) for this function:
f(x)=x^2-2x+5
----------
f(x+h) = (x+h)^2 - 2(x+h) + 5
f(x+h) = x^2 +2hx + h^2 -2x -2h + 5
---
f(x+h) = x^2 +(2h-2)x + h^2-2h+5
===========================
Cheers,
Stan H.
==================